OpenStudy (anonymous):

Please helppppppp triple sad face!!!! Elimination method. 5x + 2y = 55 2x - y = 19

OpenStudy (anonymous):

Ok for this, you basically want to get either the x or y alone in the end so you can solve. To make things simple, we will solve for x. This means that we want the y's to cancel out. 5x + 2y = 55 2x - y = 19 multiply 2x - y = 19 by 2 so the y's will cancel out 2x(2) - (2)y = 19(2) this gives us 4x -2y = 38 then we simply add the two equations together! 5x + 2y = 55 + 4x -2y = 38 ___________________ 9x = 93 divide by 9, and viola! your answer is x = 10 1/3 to get the y, you just put 10 1/3 in for any x and solve for the y

OpenStudy (anonymous):

thats not it